### v1

# Time-frequency Entangled Photon Mediated CCZ Gate

Abstract
High-fidelity native multi-qubit operations are crucial to efficient quantum circuit compilation due to their ability of shortening circuit depth and enhence the performance. However, the design and implementation of these gates remain a challenge. Here, we demonstrate a hardware-efficient scalable scheme for direct CCZ gate implementation based on two-photon absorption phenomenon, which is applicable to current superconducting quantumcomputing platforms. By carefully optimizing the parameters of qubits and couplers, we achieve a simulated fidelity over 99% within 194ns, surpassing the decomposed methods with single-qubit and two-qubit gates in both latency and overall fidelity. Crucially, the scheme is robust against parameter drifts and can be extended to CCPhase($\theta$) gates with arbitrary angles and multi-qubit operations. All these results highlight the advantages of our scheme which paves the way for substantial depth compression of complex quantum circuits for practical application in transformative quantum algorithms and simulations.
I Introduction
Despite significant advances in superconducting quantum computing over the past few decades [1, 2, 3, 4, 5, 6, 7], breakthroughs in practical quantum algorithms remain constrained by prohibitive circuit depths [8, 9, 10, 11, 12, 13]. Although single-qubit and two-qubit gates form a universal set [14], they require substantially deeper circuits than direct synthesis of native multi-qubit gates [15, 16, 17]. The Controlled-Controlled-Z $(\text{CCZ})$ gate, a pivotal three-qubit operation, exemplifies this trade-off through its decomposition into elementary gates, which introduces substantial depth overhead and non-negligible fidelity degradation relative to direct implementation. This gate imparts a $\pi$ phase shift to targeted computational basis states—a property critical for Quantum Approximate Optimization Algorithm (QAOA[18], Grover’s search [16, 19, 16], and quantum error-correcting codes [20, 21, 22]. Consequently, realizing high-fidelity native CCZ gates on superconducting processors constitutes a pivotal challenge for advancing practical quantum computing.
While the CCZ gate plays a crucial role in quantum algorithms, implementation of the high-fidelity CCZ gate remains challenging across leading quantum platforms—including superconducting circuits [23], trapped ions [24], photonic networks [25], and cavity-QED systems [26]. Existing approaches for efficient $\mathrm{CCZ}$ gates primarily adopt two paradigms: platform-dependent hardware modifications [27, 28, 29, 30] (e.g., customized couplers or quantum bus architectures) or complex optimal control pulses [31]. Both strategies incur significant experimental complexities, such as increased calibration overhead, sensitivity to parameter drift, and limited scalability—thus constraining their practical utility. Consequently, developing hardware-efficient protocols enabling direct CCZ gate implementation emerges as an essential milestone for realizing fault-tolerant quantum computation.
The phenomenon of two-photon absorption enables native three-qubit gate implementations. Initially predicted by Goeppert-Mayer in 1931 [32], this process drives substantial research interest due to applications spanning micro-nano fabrication [33, 34, 35], fluorescence imaging [36, 37], and quantum light sources [38, 39]. Two-photon absorption arises from sequential photon-induced quantum transitions. Notably, time-frequency entangled photon pairs and squeezed states enhance the absorption efficiency [40, 41, 42, 43, 44], with superconducting circuits exhibiting intrinsic capability to generate such entangled photons. Experiments have confirmed that two-photon absorption in superconducting systems induces three- and five-body interactions [45, 46, 47], thus enabling multi-qubit gates through this mechanism.
[…]
This paper is organized as follows: Section II establishes the physical model and Hamiltonian design for CCZ gate implementation in superconducting circuits. In Sec. III, we detail the gate’s implementation mechanism based on non-adiabatic geometric phases and active dynamical compensation. Section IV quantifies gate fidelity under realistic noise and demonstrates robustness against parameter drift via comprehensive numerical simulations. Section V establishes scalability pathways, extending the scheme to continuous families of three-qubit gates and multi (five and beyond)-qubit gate. Finally, Sec. VI provides concluding perspectives and discusses potential applications and future improvements.
Figure 1: (a) Toffoli gate implementations: Conventional decomposition into single-qubit and six CZ gates (left) vs. direct implementation via a native three-qubit gate (right), achieving significant gate-count reduction. (b) Schematic of the superconducting quantum circuit: Three transmon qubits coupled via two tunable couplers. Each qubit is equipped with dedicated XY and Z control lines, each coupler with an independent Z control line, and individual readout resonators facilitate simultaneous state measurement. (c) Energy-level diagram of the three-qubit system after Schrieffer-Wolff transformation (SWT)
II physical model of the system
II.1 CCZ gate
The three-qubit CCZ gate constitutes a fundamental quantum operation. The matrix representation of the CCZ gate in the two-excitation subspace basis of the three-qubit system is given by:
[…]
which applies a conditional $\pi$ phase shift exclusively when all three control qubits are in the $|111\rangle$ state as shown in Eq. (1). As a diagonal quantum operation, it preserves the computational basis while enabling the generation of three-qubit entanglement. Importantly, the CCZ gate constitutes a universal gate set for quantum computation when combined with single-qubit gates [50]—analogous to the standard $\{\mathrm{H,CZ}\}$ compilation paradigm. This universal expressivity offers a key practical advantage: Circuits compilation natively using CCZ gates exhibit significantly reduced depth compared to those decomposed into Clifford and T sequences, substantially improving their feasibility on noisy intermediate-scale quantum (NISQ) processors.
The physical realization of the high-fidelity CCZ gate represents a critical advance in quantum computation and simulation, owing to its compelling operational benefits. Such a native CCZ gate directly substitutes the standard Toffoli decomposition, reducing gate counts by approximately 89%, see Fig. 1(a). The efficiency gain originates from the gate’s innate ability to compress circuit depth beyond conventional Clifford+T compilation—paves the way for practical algorithms on NISQ-era processors, whose operation is fundamentally constrained by limited coherence times.
II.2 Physical Model
In this subsection we detail the proposed scheme. As shown in Fig. 1(b), the system consists of three adjacent transmon qubits ($\mathrm{Qubit1,Qubit2,Qubit3}$) coupled via tunable couplers ($\mathrm{Coupler1,Coupler2}$). Modeling both qubits and couplers as Duffing oscillators, the full system Hamiltonian is given by Eq. (2):
$H=H_{\mathrm{qubit}}+H_{\mathrm{coupler}}+V_{\mathrm{qq}}+V_{\mathrm{qc}},$ (2)
$H_{\mathrm{qubit}}+H_{\mathrm{coupler}}=\sum_{i}(\omega_{i}a^{\dagger}_{i}a_{i}+\frac{\alpha_{i}}{2}a^{\dagger}_{i}a^{\dagger}_{i}a_{i}a_{i}),$
$V_{\mathrm{qq}}+V_{\mathrm{qc}}=\sum_{ij}g_{ij}(a_{i}-a^{\dagger}_{i})(a_{j}-a^{\dagger}_{j}),$
where $a_{i,j}$, $a^{\dagger}_{i,j}$ $({i,j}\in\{1,2,3,c_{1},c_{2}\})$ denote the annihilation and creation operators, $\alpha_{i}$ is the anharmonicity, and $g_{ij}$ represents the direct coupling strength between elements $i$,$j$. In our proposed architecture, the frequencies of the transmon qubits can be dynamically tuned via modulating the external magnetic flux threading through their asymmetric superconducting quantum interference device (SQUID) loops [52]. The net qubit-qubit coupling can be selectively activated or suppressed through interference between direct and coupler-mediated interactions by adjusting the coupler frequencies, as demonstrated in [53]. The resulting effective coupling strengths are denoted by $\tilde{g}_{12},\tilde{g}_{23}$. To simplify the analysis, we apply the Schrieffer–Wolff Transformation (SWT[54] to decouple the couplers and incorporate high-order perturbation terms, yielding the effective Hamiltonian:
$H_{\mathrm{eff}}=\sum_{i}\frac{1}{2}\tilde{\omega}_{i}a^{\dagger}_{i}a_{i}+\tilde{g}_{12}(a^{\dagger}_{1}a_{2}+a_{1}a^{\dagger}_{2})+\tilde{g}_{23}(a^{\dagger}_{2}a_{3}+a_{2}a^{\dagger}_{3}).$ (3)
Notably, the effective coupling strengths $\tilde{g_{12}},\tilde{g_{23}}$ are explicit functions of the detuning between the computational qubits and their mediating couplers. This setup allows continuous, in-situ adjustment of the inter-qubit coupling strength by varying the coupler frequency.
Figure 2: (a) Circuit schematics for direct CCZ gate implementation, depicting the two-stage pulse sequence: three-qubits geometric phase accumulation and active dynamical phase cancellation. (b) Two quantum processes ($U_{1}$ and $U_{2}$) exhibiting destructive interference: Identically colored paths correspond to transitions induced by photons of same frequency arriving first.
III direct implementation of CCZ gate
This section presents the physical implementation of the CCZ gate, which proceeds in two complementary stages (Fig. 2(a)):
(1) Three-qubit geometric phase accumulation: A genuine three-qubits entanglement operation is realized via nonadiabatic transitions between $|101\rangle$ and $|020\rangle$ states during the first stage. A key feature is the cyclic evolution back to the computational basis state $|101\rangle$, which accumulates a geometric $\pm\pi$ phase through a closed trajectory in Hilbert space.
(2) Active dynamical phase cancellation: The dynamical phases arising from residual ZZ or ZZZ interactions during the first stage are actively canceled in this stage. This correction stage ensures the overall operation converges to a high-fidelity CCZ gate up to single-qubit phases.
Now we proceed to delineate the mechanistic details of each implementation stage.
Figure 3: (a) Schematic of the CCZ gate implementation principle: The left panel shows illustration of energy level relevant to the CCZ gate implementation. The energy splitting is tunable by adjusting $\Delta_{|101\rangle\leftrightarrow|020\rangle}$. Inset: anti-crossing between $|101\rangle$ and $|020\rangle$ while the right panel depicts utilized states (lower levels) and fidelity-limiting leakage states (upper levels), which will be analyzed in the Appendix. B. (b) depicts the two-photon absorption process with time-frequency entangled photons enabling controlled three-qubit interactions.
Figure 4: Calibration procedure. (a)(c) Calibration of transverse qubit-qubit coupling strength versus flux-bias voltage (V) in the coupler Z line. Top: Circuit schematic illustrating control lines and coupling elements. Bottom: $|101\rangle$ state population versus flux-bias voltage and time $\tau$, from which we extract the tansverse coupling strength. The dashed line marks the zero-coupling operating point. (b)(d) Calibration of the three-qubit CCZ gate. Scanning duration and amplitude of the Z pulse on $\mathrm{Qubit}2$, we measure the population of $|101\rangle$. And we identify th optimal operating point (marked as dashed line) where geometric phase accumulation reaches $\pi$. (e) Evolution trajectories of $|101\rangle$ and $|020\rangle$ states extracted from (d).
III.1 Three-qubit geometric phase accumulation
For clarity, we hereafter denote a photon at frequency $\omega_{2}+\alpha_{2}$ by $\omega_{2}^{\prime}$, and express the joint state of $\mathrm{Qubit}\ 1,\mathrm{Qubit}\ 2,\mathrm{Qubit}\ 3$ by $|\mathrm{Qubit}\ 1,\mathrm{Qubit}\ 2,\mathrm{Qubit}\ 3\rangle$, where each qubit is described by its three lowest energy levels: the ground state $|0\rangle$, the first excited state $|1\rangle$ and the second excited state $|2\rangle$. Under the resonant condition $\omega_{1}+\omega_{3}=\omega_{2}+\omega_{2}^{\prime}$, a coherent three-body interaction emerges, enabling deterministic entanglement among the three qubits [32]. But in conventional settings, random arrival times of $\omega_{2}$ and $\omega_{2}^{\prime}$ would lead to destructive interference between the $U_{1}$ and $U_{2}$ processes within second-order perturbation (see Fig. 3(a)). As a result, the net two-photon transition matrix element vanishes [55].
To circumvent this interference, we employ cascaded transitions in superconducting artificial atoms to generate time–frequency entangled photon pairs. This approach ensures that $\omega_{2}^{\prime}$ arrives prior to $\omega_{2}$, thereby suppressing the $U_{2}$ process. Therefore, destructive interference is circumvented and three-body entanglement is deterministically generated. Based on this controlled interaction, we can achieve a high-fidelity three-qubit CCZ gate and proceed to compute the effective coupling strength $J$ of this two-photon transition [45], with $J$ being given by
$J=\frac{k_{32}k_{12^{\prime}}}{\omega_{2}^{\prime}-\omega_{1}}+\frac{k_{12}k_{32^{\prime}}}{\omega_{2}^{\prime}-\omega_{3}},$ (4)
with $k_{ij}$ quantifying the coupling strength between $\omega_{i}$ and $\omega_{j}$. The condition $k_{ij}\ll|\omega_{i}-\omega_{j}|$ is naturally fulfilled in superconducting quantum circuits, strongly suppresses single-photon transitions and ensures the predominance of two-photon processes in the system’s evolution. This allows the derivation of an effective Hamiltonian for the three-qubit system:
$\tilde{H}_{\mathrm{eff}}=Ja_{1}a_{3}\Xi^{+}_{2}+\text{H.c.},$ (5)
where $\Xi^{+}_{2}$ is defined as $|2\rangle\langle 0|$, while $a_{1}$ and $a_{3}$ retain their previous definitions. Projecting Eq. (5) into the $\{|101\rangle,|020\rangle\}$ subspace yields:
$\tilde{H}_{|101\rangle\leftrightarrow|020\rangle}=\begin{pmatrix}\tilde{\omega_{|101\rangle}}&J\\ J&\tilde{\omega_{|020\rangle}}\\ \end{pmatrix}$. (6)
The three-body interaction Hamiltonian in Eq. (6) enables direct entanglement generation through coherent oscillations between the states $|101\rangle$ and $|020\rangle$. When prepared in the $|101\rangle$ state, under cyclic evolution governed by this Hamiltonian, the system deterministically returns to its initial state, accumulating a nontrivial $\pi$ phase on $|101\rangle$ state. This is different from the conventional CCZ phase (on $|111\rangle$) only by one single-qubit gate. To implement this operation, we calibrate parameters for $\pi$ phase accumulation via the pulse sequence in Fig. 4(b), which comprises three parts:
  1. i)
    State preparation: Initialize the system at $|101\rangle$ by applying single-qubit gates to $\mathrm{Qubit}1,\mathrm{Qubit}2,\mathrm{Qubit}3$;
  2. ii)
    Parameter scan: For $\mathrm{Qubit}2$, we perform a 2D sweep of flux bias (corresponds to $\Delta_{|101\rangle\leftrightarrow|020\rangle}$) and pulse width. As demonstrated in Fig. 4(b), we convert voltage coordinates to equivalent $\Delta_{|101\rangle\leftrightarrow|020\rangle}$ using pre-calibrated conversion maps;
  3. iii)
    Measurement: Measure the population $P(|101\rangle)$ and $P(|020\rangle)$ via readout resonators.
We extract linecuts from Fig. 4(d) at the working bias point, revealing coherent oscillations between $|101\rangle$ and $|020\rangle$ within a 147ns operation window with periodicity matching theoretical predictions. From the oscillation frequency, we directly determine the effective coupling strength. This measured dynamics exhibits quantitative agreement with predicted values across all explored flux biases as measured in Fig. 4(c), validating our geometric phase accumulation model.
III.2 Active dynamical phase cancellation
Crucially, while the $|{101}\rangle\leftrightarrow|{020}\rangle$ transition generates the target geometric phase, residual ZZ and ZZZ interactions concurrently introduce parasitic dynamical phases. We quantitatively characterize these error-inducing phase accumulations as follows:
$\tilde{V}^{\mathrm{ZZ}}_{\mathrm{qq}}$ $=$ $\sum_{ij}\zeta_{ij}a_{i}^{\dagger}a_{i}a_{j}^{\dagger}a_{j}.$ (7)
Within the computational subspace, the accumulated phase during Stage 1 is described by the unitary $U_{\mathrm{acc}}=e^{-\mathrm{i}(\varPhi_{s}+\varPhi_{m})}$, where $\varPhi_{s}$ denotes single-qubit phases and $\varPhi_{m}$ the multi-qubit entanglement phases. Active cancellation of $\varPhi_{s}$ via virtual-Z (VZ[56] gates yields the unitary operation:
$U=\begin{pmatrix}1&~&~&~&~&~&~&~\\ ~&1&~&~&~&~&~&~\\ ~&~&1&~&~&~&~&~\\ ~&~&~&e^{-\mathrm{i}\varphi_{|011\rangle}}&~&~&~&~\\ ~&~&~&~&1&~&~&~\\ ~&~&~&~&~&e^{-\mathrm{i}\varphi_{|101\rangle}}&~&~\\ ~&~&~&~&~&~&e^{-\mathrm{i}\varphi_{|110\rangle}}&~\\ ~&~&~&~&~&~&~&e^{-\mathrm{i}\varphi_{|111\rangle}}\end{pmatrix}$, (8)
with $\varphi_{|011\rangle}=\varphi_{23},\varphi_{|101\rangle}=\varphi_{13}+\varphi_{\mathrm{geometric}},\varphi_{|110\rangle}=\varphi_{12},\varphi_{|111\rangle}=\varphi_{12}+\varphi_{23}+\varphi_{13}+\varphi_{123}$ where $\varphi_{ij}=\int_{0}^{\tau}\zeta_{ij}\mathrm{d}t$ and $\varphi_{123}=\int_{0}^{\tau}\zeta_{123}\mathrm{d}t$ represent two-qubit and three-qubit dynamical phases, respectively. And $\varphi_{\mathrm{geometric}}$ represents the geometric phase accumulated by the nonadiabatic process. Our scheme targets exclusive accumulation of the desired phase $\pi$ on the $|101\rangle$ state while actively suppressing all other parasitic phase contributions.
Targeted CPhase gates are applied to cancel the residual ZZ and ZZZ couplings (as will be detailed in the Appendix. A). Through Ramsey-type interferometry measurements, we characterized and quantified the unwanted dynamical phases. Utilizing these target phase values, we then actively cancel them by applying non-adiabatic controlled-phase (CPhase) gates. Simulation results indicate that the required gate durations for these two cancellation pulses are $\tau_{12}$=23.4ns,and $\tau_{23}$=23.5ns respectively. We perform precision calibrations to enforce $\varphi_{13}+\varphi_{\mathrm{geometric}}=\pm\pi$. Given that $\varphi_{13}$ is a sufficiently small perturbation term in experimental realizations, we derive the approximation:
$U\approx\begin{pmatrix}1&~&~&~&~&~&~&~\\ ~&1&~&~&~&~&~&~\\ ~&~&1&~&~&~&~&~\\ ~&~&~&1&~&~&~&~\\ ~&~&~&~&1&~&~&~\\ ~&~&~&~&~&-1&~&~\\ ~&~&~&~&~&~&1&~\\ ~&~&~&~&~&~&~&1\end{pmatrix}$. (9)
Evidently, the targeted operation aligns with the CCZ gate in our protocol. Through our two-stage protocol, we can achieve a high-fidelity three-qubit CCZ gate.
Figure 5: Unitary matrix representation of the realized three-qubit CCZ gate within the computational subspace. Top: Magnitude of matrix elements. Bottom left: Real part of the unitary. Bottom right: Imaginary part of the unitary.
IV Gate performance
Figure 6: Robustness against experimental parameter drift: (a) scanning transverse coupling strength $g$ validates resilience to coupling strength fluctuation; (b) Detuning($\Delta_{|101\rangle\leftrightarrow|020\rangle}$) scan confirms robustness against frequency fluctuation.Pink curves denote simulated data and light-blue dashed lines indicate theoretical fits across both panels.
To evaluate the performance of the directly implemented CCZ gate based on our scheme, we quantify its fidelity through the following benchmark protocol [57]:
$\mathcal{F}_{\mathrm{CCZ}}=\frac{|\mathrm{Tr}({U}_{\mathrm{ideal}}{U}_{\mathrm{real}})|^{2}+\mathrm{Tr}({U}^{\dagger}_{\mathrm{real}}{U}_{\mathrm{real}})}{d(d+1)},$ (10)
where $d$ denotes the Hilbert space dimension of the quantum system ($d=8$ for three qubits system), ${U}_{\mathrm{ideal}}$ represents the unitary operation corresponding to the ideal CCZ gate, and ${U}_{\mathrm{real}}$ denotes the realized operation simulated via the Lindblad master equation. We demonstrate a gate fidelity of $99\%$ within a 194ns operation window(see Fig. 5).
Furthermore, we characterize the robustness of our CCZ gate against experimental parameter fluctuations. In real-world quantum computers, systematic drifts in control electronics and uncharacterized environmental noise induce deviations in calibrated parameters. Quantifying fidelity resilience under such perturbations is therefore essential for assessing practical viability. We model systematic uncertainties in the superconducting circuit as uncertainties in the qubit-qubit coupling strength and the detuning between $|101\rangle$ and $|020\rangle$, that is, $g\rightarrow(1+\delta)g$, $\Delta_{|101\rangle\leftrightarrow|020\rangle}\rightarrow\Delta_{working}+\zeta$, with $\delta\in[-0.05,0.05],\zeta\in[-1\mathrm{Mhz},1\mathrm{Mhz}]$.
As evidenced in Fig. 6, our directly implemented CCZ gate maintains a fidelity exceeding 92% under significant parametric deviations within operation window. This robust resilience to systematic perturbations validates our scheme’s viability for nowadays quantum processors.
Additionally, we note the presence of unexpected minor dips during parameter drift scans, which we attribute to resonant coupling between computational states and spurious modes in the system. A detailed investigation of these features will be pursued in future work.
V Extensions of this scheme
Figure 7: (a) Excited-state population of $|101\rangle\leftrightarrow|020\rangle$ transitions under parametric $\Delta_{|101\rangle\leftrightarrow|020\rangle}$ modulation. We perform a two-dimensional sweep of flux-bias amplitude and pulse duration,with golden star indicating maximal population recovery points. (b) Conditional geometric phase.The golden star corresponds to the CCphase($\pi$) gate. The inset shows the scheme of our CCPhase($\theta$) implementation.
This section delineates two extensions of our protocol: Continuous set of Three-qubit Gate Implementation and Multi (five and beyond)-qubit gate Extension. Now, we provide detailed expositions for both extensions.
V.1 Continuous set of Three-qubit Gate Implementation
The $\mathrm{CCPhase}(\theta)$ gate is a unitary operation that applies a conditional phase shift of $\theta$ to the three-qubit subspace. Its matrix representation is given by:
$U\approx\begin{pmatrix}1&~&~&~&~&~&~&~\\ ~&1&~&~&~&~&~&~\\ ~&~&1&~&~&~&~&~\\ ~&~&~&1&~&~&~&~\\ ~&~&~&~&1&~&~&~\\ ~&~&~&~&~&e^{-\mathrm{i}\phi}&~&~\\ ~&~&~&~&~&~&1&~\\ ~&~&~&~&~&~&~&1\end{pmatrix}$. (11)
This unitary operation is realized by utilizing a near-resonant interaction between the $|101\rangle$ and $|020\rangle$ states, characterized by a frequency detuning $|101\rangle$ and $|020\rangle$. The resulting exchange-type interaction induces coherent population oscillations between these states. When the interaction duration equals to a full oscillation period, the population initialized in $|101\rangleundergoes$ a complete cycle via $|020\rangle$ and returns to $|101\rangle$, acquiring a conditional phase $\phi$ in the process. This phase depends on both the detuning $\Delta_{|101\rangle\leftrightarrow|020\rangle}$ and the effective coupling strength (which is itself dependent on the detuning $\Delta_{|101\rangle\leftrightarrow|020\rangle}$). Therefore, the overall accumulated conditional phase is functionally governed by $\Delta_{|101\rangle\leftrightarrow|020\rangle}$. Through numerical simulation, we derive the explicit relationship between the detuning and the accumulated conditional phase.
As Fig. 7 demonstrates, the accumulated geometric phase during CCZ implementation can be deliberately modulated by controlling the detuning parameter $\Delta_{|101\rangle\leftrightarrow|020\rangle}$, which aligns with our analysis. This controlled phase accumulation enables a continuous set of three-qubit gates, extending the two-qubit $\mathrm{CPhase(\theta)}$ to parametrically programmable $\mathrm{CCPhase(\theta)}$ operations.
V.2 Multi(five and beyond)-qubit Gate Extension
The proposed scheme for native CCZ gate implementation can be systematically extended to multi-qubit controlled-phase gates, including five-qubit CCCCZ and higher-order operations. This scalability is achieved through a star-topology architecture where a central qubit interacts symmetrically with multiple peripheral qubits. In this configuration, resonant coupling occurs when the transition frequency of the central qubit matches the collective sum of the frequencies of the surrounding qubits, enabling efficient multi-qubit entanglement.
[…]
However, both methods exhibit distinct engineering trade-offs: the first method maintains compatibility with standard transmon frequencies without the need to redesign the sample. But it incurs exponential coherence decay at higher excitation states, which fundamentally limits fidelity through enhanced relaxation errors. And the latter preserves coherence by utilizing lower energy states but requires high-frequency operation, necessitating customized fabrication with heightened sensitivity to $E_{j}$ variations.
VI conclusion
In conclusion, we propose a scheme to directly implement a CCZ gate on a superconducting circuit, achieving a simulated fidelity exceeding 99% within a gate time of 194 ns. The approach demonstrates exceptional robustness against experimental parameter drift and substantially reduced decoherence sensitivity, owing to its ultrashort duration that remains well below typical coherence times. Furthermore, the protocol incorporates a streamlined two-step calibration process that significantly reduces the deployment complexity on physical quantum processors, facilitating experimental realization and scalability. However, due to constraints of our available hardware setup, the experimental implementation of the CCZ gate is not realized in this study, forming a key objective for our subsequent work.
Notably, the proposed architecture generalizes beyond discrete-variable gate operations by supporting continuous parametric families of three-qubit phase gates, such as the CCPhase($\theta$) gate, as well as scalable multi-qubit entangling modules that can be extended to five or more qubits. This flexibility enables fine-grained control over quantum operations and offers a versatile toolbox for circuit compilation. By significantly reducing circuit depth through native multi-qubit interactions and enhancing operational fidelity via optimized control sequences, the scheme establishes a practical and scalable pathway toward realizing near-term quantum applications in the NISQ era.
VII Acknowledgement
We would like to thank Wen-Hui Huang and Zi-Yu Tao from the Shenzhen International Quantum Academy, as well as Meng-Ru Yun from Henan Academy of Sciences, for their technical support to this work. This work was supported by the National Key R&D Program of China (No. 2024YFB4504101), the National Key R&\&D Program of China (No. 2024YFB4504600) and the National Natural Science Foundation of China (No. 62572482).
Appendix A Calibration Procedures and Numerical Validation
Here, we provide additional details of our scheme to facilitate further exploration and reproduction of our work by interested researchers. We begin by describing our experimental parameter configuration $\mathrm{Qubits\ 1,\ 2\ and\ 3}$ are configured as three identical transmon qubits, with their frequencies set to $\omega_{1}/2\pi=$ 5.018 GHz, $\omega_{2}/2\pi=$ 5.18 GHz, $\omega_{3}/2\pi=$ 4.98 GHz, respectively. The anhamonicity of each qubit is $\eta_{1,2,3}/2\pi=$ -0.35 GHz, and the effective inter-qubit coupling strength is set to $\tilde{g}_{1,2,3}/2\pi=$ 0.015 GHz.
Figure A1: Dynamical evolution of the CPhase gates employed to cancel parasitic dynamical phases in the system. The diagram traces the population evolution of the target states under the applied gates, with the vertical dashed lines marking the optimized working points at 23.4 ns (CPhase Gate 12) and 23.5 ns (CPhase Gate 23) where the accumulated conditional phase precisely counteracts the unwanted dynamical phase.
Figure A2: Quantitative evaluation of state leakage during the three-qubit CCZ gate operation. The panels from left to right show the time evolution of the leakage population to state $|110\rangle$, state $|011\rangle$, and the total leakage probability, respectively. A red pentagram in the rightmost panel highlights the operational point with minimized total leakage within the CCZgate parameter space, indicating the optimal working region for high-fidelity gate performance.
We now provide further details regarding our calibration procedure. Prior to any experimental implementation, we begin by performing individual Z-bias distortion correction for all qubits and couplers. By implementing fine-grained distortion calibration, we precisely control the qubit frequencies to their target positions—a crucial prerequisite for subsequent calibration procedures. Then each single-qubit property must be carefully calibrated. This includes the characterization of single-qubit gates and readout parameters to ensure optimal performance of each qubit at its idle point. Additionally, before proceeding with the three-qubit phase calibration, we must characterize the properties of the tunable couplers. As shown in Fig. 4(a) in the main text, this step aims to establish the relationship between the effective coupling strength $\tilde{g}$ and the applied control voltage $V$.
We now proceed to calibrate the optimal parameters for the three-qubit CCZ interaction. This is achieved by fixing the frequencies of $\mathrm{Qubit}\ 1$ and $\mathrm{Qubit}\ 3$, along with the voltages applied to the $\mathrm{Coupler}\ 1$ and $\mathrm{Coupler}\ 2$, while scanning both the frequency of $\mathrm{Qubit}\ 2$ and the interaction time. The system’s state is subsequently measured to identify the optimal operational regime. The optimal operational point in this step can be readily identified from the two-dimensional scan—simply locating the star-marked point in Fig. 4(d) of the main text. A more critical task, however, is the accurate characterization of the dynamical phases accumulated during this process, which must be actively canceled in the subsequent calibration stage. As shown in a previous study [48], these phases can be measured through initializing three qubits in six special states, i.e., $|0\!+\!0\rangle$$|1\!+\!1\rangle$$\rangle|0+1|0\!+\!1\rangle$|1 0+|1\,0\ $\rangle$, and |0 0+|0\,0\ $\rangle$. We then measure all conditional phases through Ramsy-Like experiments, i.e., $\varphi_{12}=\varphi_{|1+0\rangle}-\varphi_{|0+0\rangle}$, $\varphi_{23}=\varphi_{|0+1\rangle}-\varphi_{|0+0\rangle}$, $\varphi_{13}=\varphi_{|1\,0\,\!+\!\rangle}-\varphi_{|0\,0\,\!+\!\rangle}$, $\varphi_{123}=\varphi_{|1+1\rangle}-\varphi_{|0+0\rangle}$, where $\varphi_{12},\varphi_{23},\varphi_{123}$ are defined in the main text. With these characterized parameters, we can eliminate the residual phases via targeted CPhase gate operations, thereby obtaining a purified CCZ gate.
Appendix B Gate Leakage and Error Mechanism Analysis
In this section, we present a systematic analysis of state leakage and error sources in our implementation of the CCZ gate. Particular attention is given to the unintended population transfer to the leakage states $|011\rangle$ and $|110\rangle$, in addition to the intended computational subspace $|020\rangle$ and $|101\rangle$. We quantify the leakage error, identify its physical origins, and evaluate its impact on gate performance. We model the system using the full Hamiltonian including all relevant interactions. The leakage to states $|011\rangle$ and $|110\rangle$ is primarily attributed to frequency crowding during CCZ operation, resulting in parasitic crosstalk. The leakage rate to each state is quantified through numerical integration of the Lindblad master equation:
$\mathcal{L}_{|011\rangle}=\operatorname{Tr}\left[|011\rangle\langle 011|\cdot\rho(t_{\text{CCZ}})\right],$ (12)
$\mathcal{L}_{|110\rangle}=\operatorname{Tr}\left[|110\rangle\langle 110|\cdot\rho(t_{\text{CCZ}})\right],$ (13)
where $\rho(t_{CCZ})$ represents the density matrix at the end of the operation.The total leakage error is calculated as :
$\mathcal{L}_{total}=\mathcal{L}_{|011\rangle}+\mathcal{L}_{|110\rangle}.$ (14)
We identify a point near the CCZ gate operating point that minimizes leakage error, as shown in Fig. A2. The data in Fig. A2 indicates the existence of an operational window within the parameter region of interest where leakage error is minimized. Identifying and operating at this point is critical for minimizing the deleterious effect of leakage on the overall gate fidelity. Within a gate duration of 147 ns, the measured leakage population to state $|011\rangle$ and $|110\rangle$ remains as low as 2.5%. Moreover, at this specific point, the accumulated three-qubit phase is closest to $\pi$, making it the optimal operating point for our CCZ gate implementation.
Several strategies can be employed to mitigate such leakage errors. As noted above, one primary cause of this leakage is the close proximity of their frequency transitions within the operational parameter range. Our study reveals that the energy level separation between these states depends on the magnitude of the anharmonicity $\eta_{1,2,3}$. A larger absolute value of $\eta_{1,2,3}$ results in a greater energy level difference, thereby reducing leakage error. Consequently, one could intentionally design superconducting qubits with higher charging energy $\mathrm{E_{c}}$ (and thus larger anharmonicity) to implement this scheme with improved performance. Although this approach represents a promising direction, our work focuses on compatibility with mainstream quantum platforms; therefore, further exploration of such device-level modifications is beyond the scope of this study.

### v2

@@
-Time-frequency Entangled Photon Mediated CCZ Gate
+revtex4-2Repair the float
 Abstract
@@
 This gate imparts a $\pi$ phase shift to targeted computational basis states—a property critical for Quantum Approximate Optimization Algorithm (QAOA[18], Grover’s search [16, 19, 16], and quantum error-correcting codes [20, 21, 22].
-Consequently, realizing high-fidelity native CCZ gates on superconducting processors constitutes a pivotal challenge for advancing practical quantum computing.
+Consequently, realizing high-fidelity native CCZ gates on superconducting processors is critical for advancing practical quantum computing.
 While the CCZ gate plays a crucial role in quantum algorithms, implementation of the high-fidelity CCZ gate remains challenging across leading quantum platforms—including superconducting circuits [23], trapped ions [24], photonic networks [25], and cavity-QED systems [26].
@@
 Each qubit is equipped with dedicated XY and Z control lines, each coupler with an independent Z control line, and individual readout resonators facilitate simultaneous state measurement.
-(c) Energy-level diagram of the three-qubit system after Schrieffer-Wolff transformation (SWT)
+(c) Energy-level diagram of the three-qubit system after Schrieffer-Wolff transformation (SWT).
 II physical model of the system
@@
 (1) Three-qubit geometric phase accumulation: A genuine three-qubits entanglement operation is realized via nonadiabatic transitions between $|101\rangle$ and $|020\rangle$ states during the first stage.
-A key feature is the cyclic evolution back to the computational basis state $|101\rangle$, which accumulates a geometric $\pm\pi$ phase through a closed trajectory in Hilbert space.
+A key feature is the cyclic evolution back to the computational basis state $|101\rangle$, which accumulates a geometric $\pi$ phase through a closed trajectory in Hilbert space.
 (2) Active dynamical phase cancellation: The dynamical phases arising from residual ZZ or ZZZ interactions during the first stage are actively canceled in this stage.
@@
 Scanning duration and amplitude of the Z pulse on $\mathrm{Qubit}2$, we measure the population of $|101\rangle$.
-And we identify th optimal operating point (marked as dashed line) where geometric phase accumulation reaches $\pi$.
+And we identify th optimal operating point (marked by a red pentagram on the dashed line) where geometric phase accumulation reaches $\pi$.
 (e) Evolution trajectories of $|101\rangle$ and $|020\rangle$ states extracted from (d).
@@
 But in conventional settings, random arrival times of $\omega_{2}$ and $\omega_{2}^{\prime}$ would lead to destructive interference between the $U_{1}$ and $U_{2}$ processes within second-order perturbation (see Fig.
-3(a)).
+2(b)).
 As a result, the net two-photon transition matrix element vanishes [55].
 To circumvent this interference, we employ cascaded transitions in superconducting artificial atoms to generate time–frequency entangled photon pairs.
-This approach ensures that $\omega_{2}^{\prime}$ arrives prior to $\omega_{2}$, thereby suppressing the $U_{2}$ process.
+In Fig.
+3, we illustrate the physical realization of this process using superconducting qubits, which is achieved by resonantly coupling the $|101\rangle$ and $|020\rangle$ states.
+This facilitates an energy exchange in which Qubit 2 loses two photons while Qubit 1 and Qubit 3 each absorb one, even though each absorption process is individually off-resonance.
+By facilitating this, the approach ensures that $\omega_{2}^{\prime}$ arrives prior to $\omega_{2}$, thereby suppressing the $U_{2}$ process.
 Therefore, destructive interference is circumvented and three-body entanglement is deterministically generated.
@@
 $J=\frac{k_{32}k_{12^{\prime}}}{\omega_{2}^{\prime}-\omega_{1}}+\frac{k_{12}k_{32^{\prime}}}{\omega_{2}^{\prime}-\omega_{3}},$ (4)
-with $k_{ij}$ quantifying the coupling strength between $\omega_{i}$ and $\omega_{j}$.
+where $k_{ij}$ quantifying the coupling strength between $\omega_{i}$ and $\omega_{j}$.
 The condition $k_{ij}\ll|\omega_{i}-\omega_{j}|$ is naturally fulfilled in superconducting quantum circuits, strongly suppresses single-photon transitions and ensures the predominance of two-photon processes in the system’s evolution.
@@
 Simulation results indicate that the required gate durations for these two cancellation pulses are $\tau_{12}$ = 23.4 ns, and $\tau_{23}$ = 23.5 ns respectively.
-We perform precision calibrations to enforce $\varphi_{13}+\varphi_{\mathrm{geometric}}=\pm\pi$.
+We perform precision calibrations to enforce $\varphi_{13}+\varphi_{\mathrm{geometric}}=\pi$.
 Given that $\varphi_{13}$ is a sufficiently small perturbation term in experimental realizations, we derive the approximation:
@@
 In conclusion, we propose a scheme to directly implement a CCZ gate on a superconducting circuit, achieving a simulated fidelity exceeding 99% within a gate time of 194 ns.
-The approach demonstrates exceptional robustness against experimental parameter drift and substantially reduced decoherence sensitivity, owing to its ultrashort duration that remains well below typical coherence times.
+The approach demonstrates exceptional robustness against experimental parameter drift and substantially reduced decoherence sensitivity, owing to its short duration that remains well below typical coherence times.
 Furthermore, the protocol incorporates a streamlined two-step calibration process that significantly reduces the deployment complexity on physical quantum processors, facilitating experimental realization and scalability.
@@
 4(a) in the main text, this step aims to establish the relationship between the effective coupling strength $\tilde{g}$ and the applied control voltage $V$.
-We now proceed to calibrate the optimal parameters for the three-qubit CCZ interaction.
+Subsequently, we describe the procedure for determining the optimal parameters required to realize the three-qubit CCZ interaction.
 This is achieved by fixing the frequencies of $\mathrm{Qubit}\ 1$ and $\mathrm{Qubit}\ 3$, along with the voltages applied to the $\mathrm{Coupler}\ 1$ and $\mathrm{Coupler}\ 2$, while scanning both the frequency of $\mathrm{Qubit}\ 2$ and the interaction time.